A large rock is released from rest from the top of a tall building. The average speed of the rock during the first second of the fall is 5 m/s. What is the average speed of the rock during the next second? (In this question we use the approximate value of 10 m/s2 for the gravitational acceleration.)
Answer #1

Let us assume the downward direction as positive.

Gravitational acceleration = g = 10 m/s2

Initial velocity of the rock = V0 = 0 m/s (Released from rest)

Velocity of the rock at the end of the first second = V1

Time period = T1 = 1 sec

V1 = V0 + gT1

V1 = 0 + (10)(1)

V1 = 10 m/s

Distance traveled by the rock in the 2nd second = D2

Time interval = t = 2 - 1 = 1 sec

D2 = V1t + gt2/2

D2 = (10)(1) + (10)(1)2/2

D2 = 15 m

Average speed of the rock during the 2nd second = V2avg

V2avg = 15 m/s

Average speed of the rock during the 2nd second = 15 m/s